An experimenter flips a coin 100 times and gets 62 heads. We wish to test the claim that the coin is fair (i.e. a coin is fair if a heads shows up 50% of the time). Test if the coin is fair or unfair at a 0.05 level of significance. Calculate the z test statistic for this study. Enter as a number, round to 2 decimal places.
Answer #1

H0: p = 0.50

Ha: p 0.50

Sample proportion = 62 / 100 = 0.62

Test statistics

z = ( - p ) / sqrt [ p (1 - p) / n ]

= ( 0.62 - 0.50) / sqrt ( 0.50 * 0.50 / 100 )

= 2.40